How do you solve 100(X+1/2)^2=1
Answer by Alan3354(69443)

How do you solve 100(X+1/2)^2=1
------------
Divide by 100
(X+1/2)^2 = 1/100
Take sqrt of both sided
x + 1/2 = 1/10 or -1/10
x = 1/10 - 1/2
x = -2/5
----------
x = -1/10 - 1/2
x = -3/5